OpenStudy (anonymous):

sin theta-sin theta*cos^2 theta=sin^3 theta

OpenStudy (sidsiddhartha):

u wanna prove it?

OpenStudy (sidsiddhartha):

@xcandacex3

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

yeah take sin theta common in the left hand side \[\sin \theta-\sin \theta*\cos^2 \theta=\sin \theta(1-\cos^2 \theta)=\sin \theta*\sin^2\theta=\sin^3 \theta\]

OpenStudy (sidsiddhartha):

use this: \[\sin^2 \theta+\cos^2 \theta=1\\1-\cos^2 \theta=\sin^2 \theta\]

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

yes THANK YOU!! :)

OpenStudy (sidsiddhartha):

yw!!